OpenStudy (idealist10):

Solve (3ycosx+4xe^x+2x^2*e^x)dx+(3sinx+3)dy=0.

OpenStudy (idealist10):

@amistre64 @hartnn @dumbcow

OpenStudy (idealist10):

Please help me. This is an exact differential equation.

OpenStudy (amistre64):

if its exact then integrate one side, then derive it to the other side, is what i recolect

OpenStudy (idealist10):

Can you show your work, amistre?

OpenStudy (dumbcow):

ok i dont have a lot of time to work through this, i will get you started http://tutorial.math.lamar.edu/Classes/DE/Exact.aspx \[\int\limits N dy = \int\limits (3 \sin x +3) dy = C\] \[3 y \sin x +3y +h(y) = C\]

OpenStudy (amistre64):

int: (3+3sinx) dy = 3y-3ysin(x)+f(x) Dx: 3y-3ysin(x)+f(x) = -3ycos(x)+f'(x) now compare -3ycos(x)+f'(x) = 3ycosx+4xe^x+2x^2*e^x

OpenStudy (amistre64):

the sites starting to lag ... bummer

OpenStudy (amistre64):

i see i fubarred a neative

OpenStudy (amistre64):

3ycos(x)+f'(x) = 3ycosx+4xe^x+2x^2*e^x f'(x) = 4xe^x+2x^2*e^x

OpenStudy (dumbcow):

oops i meant h(x) then differentiate with respect to x \[\rightarrow 3y \cos x +h'(x) = \frac{d}{dx}(3 \sin x +3) = 3 \cos x\] \[h'(x) = 3 \cos x (1-y)\]

OpenStudy (amistre64):